Update 1st February

Added by Kyberd, Paul about 8 years ago

Step IV:

Work in the MICE Hall is focussed on the necessary preparations for the LiH scattering measurements that are planned for the next ISIS User Cycle. It has been decided to adjust the positions of the spectrometer solenoids to account for the measured offset between the magnetic and mechanical axes. Before this work can be undertaken, appropriate bellows to join the spectrometer solenoids to the focus-coil module must be procured.

Many “house-keeping” tasks have been completed (e.g. flushing of the beam-line magnets, DKsoln fridge and compressor servicing, implementation of “pseudo-kinematic” locators for TOF2, KL and EMR …).

Work in R9 is focussed on preparing for the extraction of the liquid-hydrogen vessel so that the leaking seal can be repaired.

A new quench-detection/quench-protection system for the spectrometer solenoids has been devised by the spectrometer-solenoid team. Detailing of the specification of the various components is under way and measurements of critical parameters will be made in Rack Room 2 this week. The spectrometer-solenoid recovery plan continues to be developed.

Software and analysis:
The systematic check-out of the tracker software (Dobbs, Heidt, Hunt, Overton, Uchida) is reaching the end of the first pass through the code. The sign-flip issues in the reconstruction have been addressed and the algorithm check-out from cluster finding, through pattern recognition to the track fit has been completed successfully. Features in the tracker simulation have also been understood and addressed. A. Dobbs is preparing a new release of MAUS that will contain fixes to all known issues. With this release, D. Rajaram will make a consistent reprocessing of all the data that was taken in 2015.

When the data is reprocessed, the alignment constants derived by F. Drielsma will be included. The first link shows the status of the alignment of the instrumentation from TOF1, through both the trackers to TOF2, KL and EMR.

The analysis of the data taken on the 7th October 2015 with the end and centre coils of the upstream solenoid at nominal current is progressing. The second link (Blackmore) shows the emittance reconstructed from tracks in the upstream tracker plotted in 5 MeV/c momentum bins from 160 MeV/c to 245 MeV/c. The data has been split into two samples of equal size. The first is used to calculate the emittance at the entrance to the upstream tracker (station 5, black solid circles). The second is used to calculate the emittance at the exit of the tracker (station 1, red solid triangles). While more work needs to be done, the plot indicates that we shall be able to make a good measurement of emittance using the single-particle technique.

The xenon scattering data taken just before Christmas is also being analysed. The third attachment shows the scattering angle determined by taking the dot product between the track direction determined in the upstream tracker with that determined downstream (Nugent). The scattering analysis needs to include the alignment corrections outlined above, but, the preliminary analysis indicates that it will be possible to unfold the scattering distribution.

Video conference:
The next video conference will take place on Thursday 04Feb16.

MICE update 25 July 2015

Added by Kyberd, Paul over 8 years ago

Data taking and operations:

ISIS Cycle 2015/01b ended yesterday. During the night of 24Jul15, the experiment was operated successfully with 100A in the downstream spectrometer solenoid. As you’ll have seen from Chris Rogers’ email, helix tracks were observed in the downstream tracker. In the first attachment you’ll find the event that Chris circulated.

The first helical track is the tip of the iceberg in terms of the work that it took to get here.

Over the past month there has been an enormous effort by the US magnet team. Led by Mark Palmer and Alan Bross the team were able to start the training of the downstream solenoid and to bring it into stable operation at 100A for the data taking during the night of 24Jul15. On Friday, 24Jul15, SSD achieved 190A — around two thirds of the operating current.

My congratulations, and thanks, to Maria Leonova and Sandor Feher and all of the magnet team. In particular Pierrick Hanlet has made an enormous effort to serve the controls needs of the magnet commissioning while keeping an eye on the experiment as well.

The Operations team too; ably coordinated by the MOMs Ryan Bayes and, in the last few days, Paul Hodgson has worked hard to keep the experiment going in a period when the daily changes in schedule were challenging. More over, the reconstruction, and first analysis of the data, was completed in less than 24 hours; this is an important capability and one which we will sorely need as we go forward in Step IV. My congratulations, therefore, to the Operations, S/w&C and analysis teams.

Progress has also been made on the mechanical alignment of the trackers using the field-off data collected in Cycle 2015/01a. For information I attach plots of the comparison of position and slope of the tracks between the upstream and downstream trackers prepared by Chris Hunt. There is still much to understand in the data, however, with the field-off and field-on data combined, I believe we have a good basis for studying the magnetic alignment of the experiment.

First Track


Looking forward:

We have much to look back on with satisfaction. However, the last few weeks were not without some challenges, particularly in terms of carrying out the operations necessary to commission the magnets.

Permission has been granted to operate the magnets at Step IV. To gain permission to operate, procedures and responsibilities have been discussed and agreed between MICE and ISIS. Now that Cycle 01b is complete, and first data with field has been taken, it will be necessary to review and refine the procedures to make sure they are fit for purpose and are able to guarantee safety while allowing efficient working. This process will begin first thing next week.

During the preparation for training, a fault was discovered in the upstream solenoid. The magnet is presently being warmed up to allow the spectrometer-solenoid team to inspect the warm feedthroughs where the problem is expected to be. The remedial work necessary to bring the magnet back online will be given priority with a view to starting the cool-down of the upstream solenoid in the second week of August.

Once the repair has been effected and the MICE Hall prepared for continued magnet training, the commissioning of the spectrometer solenoids will recommence and the focus-coil commissioning will begin. The timescales are such that it is likely that a portion of Cycle 2015/02 will be required for magnet training. While this is regrettable, it is now important to consolidate our procedures and our expert teams. If we can do both of these things effectively over the coming month or so, we will be in a much stronger position to execute our data-taking campaign for Step IV.

Over the coming weeks the project, operations and analysis teams need to work together to revise the operations programme for Cycle 2015/02 and beyond. 24/7 shift cover will be necessary to provide support for commissioning and operations and, of course, for data taking. So, please do respond to the request made by Paul Kyberd to volunteer for shifts in Cycle 2015/02.

And finally …

A lot has been accomplished and much valuable experience has been gained. It is now important to learn from that experience to build the platform on which we can stand to deliver Step IV.

Update 13 July

Added by Kyberd, Paul over 8 years ago

The purpose of this update is to summarise the state of the experiment at the end of ISIS Cycle 2015/01a and to outline the priorities for the coming Cycle (2015/01b), which is about to start.

Status update:

By the end of 2015/01a, all of the MICE instrumentation had been commissioned and detailed monitoring had been made available in the control room. As an example, the first attachment shows a selection of the plots available from the Online reconstruction. Each of the detector systems is now represented.

Offline, substantial progress had also been made. MAUS now is able to reconstruct the data faster that we can record it; a really important step. The straight-track reconstruction too has been debugged and refined. As an example of the progress here, the second attachment shows the smoothed residuals from the track fit in each plane of both trackers. There are details to understand, but, the data and the reconstruction software can now be used to study the alignment, efficiency and calibration of the devices.

For the past week or so, the spectrometer solenoids have been cold, each containing a volume of liquid helium. The control system has maintained the magnets at a stable operating point. To reach this state required a substantial, dedicated effort from the US magnet team. The third attachment shows the strip-chart of the cool down; you can see where the upstream and downstream solenoids reach their operating temperature.

The cool-down of the focus-coil continues to go to plan. The magnet is now at around 80K and is expected to reach liquid-helium temperature at the end of this week.

Over the past few weeks substantial progress has been made towards completing the electrical and control systems for the magnets. This work culminated at the weekend with the start of quench-detector-system tests. The hard work and close coordination of Pierrick, the US magnet team and the DL electrical and controls teams have made this possible.

Commissioning and operations in ISIS Cycle 2015/01b

The electrical and controls experts are now preparing the systems for the start of magnet training. The goal is that magnet training will start in the week of the 20Jul15. In the run up to training, many systems need to be checked out and a formal “sign-off” procedure is being developed.

Tomorrow, TOF2, KL and EMR will be moved into their final Step IV positions. This will necessitate a further round of TOF calibrations. Steve Boyd, in consultation with the Analysis group and others, has developed a run plan. Data taking must be scheduled around magnet commissioning, so, items on the run plan will be executed in priority order as time permits.

It is important for the development of the experiment that the training of the MICE magnets is expedited. Therefore, in the coming days, preparations for magnet training, and then the magnet training itself, will have priority. Since we are setting up a new system and then training the magnets together for the first time, this period will necessarily be one in which plans are fluid and subject to change. I would therefore ask the shift crews to bear with, and support, Steve Boyd and Ryan Bayes (the MOM) as they respond to the evolving needs of the experiment.

This is an exciting time for MICE; are close to bringing Step IV into operation

News July 1st

Added by Kyberd, Paul over 8 years ago

Preparation of Step IV operations:

The cool-downs of the spectrometer solenoids and the focus-coil module are now well underway. Our goal is to initiate the training of the spectrometer solenoids on Friday 03Jul15; this week! To do this we require safety-sign off from ISIS to obtain permission to operate. To obtain the necessary safety approval our documentation will be reviewed on Thursday (02Jul15) morning. This meeting will be followed by a safety tour of the MICE Hall on Thursday afternoon.

As part of our preparations for the safety tour we are making a “deep clean” of the MICE Hall during which the various pieces of material and equipment that are presently stored in the MICE Hall will be moved into storage elsewhere on site. John Govans and Craig Macwaters are masterminding this clear-up.

Given the timetable, the clear-up needs to move forward right away. Please help John, Craig and the team in the Hall by communicating any special requirements that you may have to John or Craig as soon as possible.

First processing of Step IV data to date:

We’ve taken quite a lot of data during the ISIS Cycles 2014/03 and 2015/01a. It is important that we analyse efficiently the data that has been taken to extract important messages for the coming runs. Therefore, our software and computing team are preparing a first processing of the data using an appropriate and consistent set of calibration and geometry constants. We plan to carry out this first processing before the start of ISIS Cycle 01b.

The schedule that the s/w&c team seek to hold is:

• 06Jul15:

o List of runs to be processes prepared by the MOM (Ryan Bayes);
o List of updates to the ConfigDB (geometry and calibration) prepared by the MOM;

• 08Jul15:

o Updated ConfigDB released for checking by system experts;
o Start of “test processing” of a subset of the data;

• 10Jul15:

o Results of test processing released to system experts for checking;

• 11/12Jul15:

o Start of processing of Step IV data to date.

Please give Ryan and the s/w&c team your full support when they ask for checks etc. Its important to the experiment to get this data pushed through to the analysers.

MICE Update 17 May 2015

Added by Kyberd, Paul almost 9 years ago

Dear All,

Progress update:

Step IV:

The material for the north PRY arrived at RAL on Tuesday 12May15. When work to install nine (of ten) cryocoolers and the power and sensor cables to the spectrometer solenoids had been completed on Wednesday 13May15, the north PRY legs were installed on Thursday 14May15. The legs were aligned and bolted down on Friday.

In parallel, work on the absorber/focus-coil module continued. The module is now almost ready to be placed on the beam line. The bellows that link the focus-coil module to the spectrometer solenoids have been prepared and will be installed at the start of next week. The focus-coil module will then be moved onto the beam line.

The target team have spliced and terminated the new fibres at junction boxes inside and outside the synchrotron. Work over the coming days will ensure the integrity of the new fibres so that the target may be operated with them.

Progress on LH2:

Excellent progress has been made on preparing the LH2 system for Step IV operation. The integrity of the safety windows has been proven to the satisfaction of the MICE/ISIS H2 working group. Further, the H2 exhaust path has been agreed. These are essential (necessary but not sufficient) steps towards the operation of Step IV with LH2.

The exhaust path now has to be proved and the system commissioned with He gas. In addition, the necessary documentation must be finalised to allow the MICE/ISIS Safety Committee to recommend to the Directors of ISIS and PPD that the LH2 system is fit for operation.

At present we have decided that the first absorber to be studied at Step IV will be LiH. Only once the technical, safety and formal sign-off milestones outlined above are completed can this decision be revisited.

The near future:

MICE Hall:

The big task next week is the installation of the north PRY and the trial assembly of the PRY end plates. The line will be made ready on Monday and Tuesday. Over this period the bellows will be fitted, the absorber/focus-coil module will be placed on the beam line and the channel will be bolted together. On Wednesday, the installation of the north PRY will start. It will be important to complete the installation of the PRY next week to avoid a stack-up of work towards the end of the month.

The last of the cryocoolers in the downstream solenoid will be fitted in the week of the 25th May2015. At this time a short to ground that has been observed in the current leads of the downstream solenoid will also be rectified.


The operations plan prepared by the Physics group for ISIS cycle 2015/01a (June) was accepted by the EB at its meeting on 08May15. The Operations group is now working on implementing the plan.

Operations in June and July will be devoted to bringing the spectrometer solenoids and focus-coil module up to operating current, commissioning the tracker readout with beam and performing initial alignment studies of the Step IV apparatus. In June there will be one data-taking shift per day. The shift will be staffed by two shifters and run from 22:00 to 06:00. A “floating shifter” will be in place for the day time (06:00 to 22:00) to support experts in commissioning. Over this period, priority will be given to magnet training and tracker-readout commissioning.

From cycle 2015/01b (July), 24/7 operations (three shifts of two shifters each) will begin. While still scheduled as a commissioning period, it will be the ambition to move more and more to a routine data-taking mode.

In the light of all this excellent progress, I think we can agree that Step IV operations can formally be scheduled to begin in ISIS cycle 2015/02 (August 2015).


1. The MICE instance of the ISIS User Database is now in production. The link to it is:
⁃ Please take a moment to register or to make sure that your existing registration is up to date.
⁃ This is important as information such as your training status is held on the database. Keeping this up to date is part of our safety case and will be scrutinised by ISIS as we seek permission to operate Step IV.
2. The WWW site of CM42 is now live. You can reach it by following the link from the MICE home page:
⁃ Please register!

⁃ With best wishes as always …

⁃ Ken

MICE Update 22 March

Added by Kyberd, Paul almost 9 years ago

Progress update:

Step IV:

The plates that will form the south side of the PRY have been collected by the shipping agent from Keller in New York. The present expectation is that the plates will arrive in the UK next Tuesday and, if all goes well, be received at RAL on Wednesday.

The long-day working in the Hall has been successful and has cleared a number of jobs. Work in the Hall has been organised so as to allow the PRY plates to be delivered directly to the MICE Hall.


Weekend running in the present ISIS User Run [2014(03)] is now underway. The goals set for running this weekend (21/22 March 2015) were to establish smooth operation from target to TOF1, carry out an initial calibration of TOF0 and TOF1, take data to allow a first test of the reconstruction algorithms required for MICE Muon Beam commissioning and the completion of BLOC training using beam on target. It is to the great credit of all those involved in the preparation of our online, computing, data-handling and operations activities that all these goals have all been achieved.

Work continues to resolve the noise and related issues in the tracker readout. The installation of the experimental ground in the MICE Hall has also progressed. The work on establishing the smooth operation of the tracker readout remains a high priority.

The pattern of weekly Operations meetings (in person 11:00 Friday’s and in-person/teleconference 15:00 Monday’s) is now established. A full debrief on the weekends running will take place at the Operations meeting on Monday 23Mar15.

The near future:

MICE Hall:

Installation of the PRY plates to the south side will start on Monday 30Mar15.

Work on the cabling, compressor-hose support structures and other infrastructure will continue in parallel.

Weekend operations:

The goals presently set for operations in the present ISIS User Run are:
  • 28/29Mar15:
    ◦ Perform calibration of TOF0 and TOF1;
    ◦ Take data to allow a comparison of the performance of the old and new trigger systems;
    ◦ Take pion-beam data to allow the reconstruction algorithms required for MICE Muon Beam commissioning to be developed.
  • 11/12Apr15, 18/19Apr15 and 25/26Apr15:
    ◦ Check the TOF calibration. It will be important to take some data to be sure that the calibration is stable;
    ◦ First pass of MICE Muon Beam (MMB) commissioning:
    ▪ Start with the currents used in Step I to check the degree of reproducibility of particle rates;
    ▪ Based on the revised simulation of the beam line (Jaroslaw Pasternak), start the search for the optimum magnet settings;
    ◦ Commissioning and calibration of TOF2, KL and EMR will begin as soon as the south-side PRY is complete and the magnets and downstream detectors are in their final Step IV positions.


The MICE WWW pages continue to be improved and, of course, Durga Rajaram is always ready to receive comments and suggestions for improvements.

One important update, for those giving talks or preparing documents, is that the “public information” page has been updated and now includes not only published figures and tables but also the sketches and plots for the cooling demonstration that we agreed to make public at the recent CM. Please use these figures.

To get to the information, go to the MICE www site ( take the drop-down menu “Documents” and select “Publications”. You can also get there by taking the drop-down menu “Resources” and selecting “Drawings and Images” and then following the link to publications and figur

MICE Update 16 March

Added by Kyberd, Paul almost 9 years ago

Progress update:

Step IV:

The south-side PRY support legs are now in place and have been surveyed. The shims have been measured and machined and will be installed this week. The Hall is therefore essentially ready for the installation of the south-side PRY plates. In New York, the south-side plates are being made ready for shipment. The teams on both sides of the water are working to ensure that shipping delays are reduced to a minimum. Work at Keller continues to complete the north-side plates. So far, the critical work to install the PRY remains on track.

The “two-shift”, or perhaps more accurately the long-day working, in the Hall has started as planned today (Monday 16Mar15). Work has now started on the prioritised list of jobs. To help make sure all are aware of what is going on, the list of jobs to be carried out over the next two weeks is posted on the screen located next to the personnel door to the MICE Hall.

Software and computing:

Durga Rajaram and the online, controls and offline software teams are pressing forward to make sure all is in place for Step IV. Automatic batch processing has been established and will be put to the test in the upcoming weekend operations.

Issues to do with the event-processing time are being addressed. Changes in the raw-data format made necessary by upgrades to the online system require that the unpacking routines be updated. The necessary upgrades are being implemented.

Work continues on the reconstruction algorithms. The detailed work necessary to ensure that the Kalman track fit in the tracker is statistically correct is making steady progress (Chris Hunt). Getting this right is, of course, central to delivering the Step IV programme.


Shift sign-up for the weekend running in ISIS User Run 2014(03) (i.e. now to end April) is pretty much complete. Sign-up for the running from June (ISIS User Run 2015(01)) is now live. After a brief flurry of activity, the sign-up for shifts in June and beyond has slowed down; please do sign-up for your shifts as soon as you are able.

The near future:

MICE Hall:

While the PRY plates are in transit, work to advance the cabling, compressor-hose support structures and other infrastructure will continue.

Weekend operations from 21Mar15:

  • The service of the Linde refrigerator serving the decay solenoid starts today. The decay solenoid will not be available until after Easter. While the plan for operations is likely to have to respond to pressures from the construction and commissioning activities, the goals presently set for operations in the present ISIS User Run are:
  • 21/22Mar15, 28/29Mar15:
    ⁃ Establish smooth operation from target to TOF1;
    ⁃ Carry out an initial calibration of TOF0 and TOF1;
    ⁃ Take data to allow a comparison of the performance of the old and new trigger systems;
    ⁃ Take pion-beam data to allow a first test of the reconstruction algorithms required for MICE Muon Beam commissioning;
    ⁃ Complete BLOC training on target tuning.
  • 11/12Apr15, 18/19Apr15 and 25/26Apr15:
    ⁃ Check the TOF calibration. It will be important to take some data to be sure that the calibration is stable;
    ⁃ First pass of MICE Muon Beam (MMB) commissioning:
    ⁃ Start with the currents used in Step I to check the degree of reproducibility of particle rates;
    ⁃ Based on the revised simulation of the beam line (Jaroslaw Pasternak), start the search for the optimum magnet settings;
    ⁃ Commissioning and calibration of TOF2, KL and EMR will begin as soon as the south-side PRY is complete and the magnets and downstream detectors are in their final Step IV positions.

Once the south-side PRY is installed and the magnets are positioned on the beam line (present milestone date 14Apr15), the tracker waveguides can be installed to the south-side VLPC cryostats. The tracker team is discussing how best to make use of the time from around the 14Apr15 to advance the tracker commissioning and to profit from the weekend running.

Operations meetings:

The second MICE/ISIS Operations Meeting will take place at 15:00 today (16Mar15). The first of the regular weekly MICE Operations meetings will take place (exceptionally!) tomorrow (17Mar15) starting at 14:00. After that the pattern of operations meetings will be as we agreed:

• Friday’s 11:00: In person Operations meeting;
• Monday’s 15:00: A meeting to debrief on the weekend’s operation and to plan the coming week.

MICE Update 9 March: Part 2

Added by Kyberd, Paul almost 9 years ago

Preparations for the April 2015 international review of the project (RLSR/MPB/FAC): 

The MICE Project Board will meet on the 16th and 17th of April 2015.  The MPB meeting will be preceded by the Resource Loaded Schedule Review of the project on the morning of the 16Apr15 and the project will be required to report to the international Funding Agency Committee on the afternoon of the 17Apr15.
The last reviews in November were important for us; they went well and the project was re-baselined.  The reviews in April will give us the chance to demonstrate our command of the Step IV construction project, to show that we’re on top of our preparations for Step IV operations and physics analysis and to present the excellent progress that has been made in defining the cooling demonstration configuration and the necessary engineering details.
To achieve all of this we’ll have to deliver the necessary paperwork and prepare an excellent set of talks.  Work on the preparation for the April review of the project will start next week.  Success in April will be based on the hard work of all; please respond in line with the necessary deadlines if you receive a request for to contribute to the preparations!

MICE Update 9 March: Part 1

Added by Kyberd, Paul almost 9 years ago

Progress update:

Step IV:

The first set PRY legs have made their progress around the Hall.  The shims were installed under the legs on the North Side and the survey was repeated.  When all were satisfied, the legs were removed and the spectrometer solenoids and the focus-coil module were moved into the “park” position.  At the end of the week, the downstream set of legs had been installed on the South Side.  Final positioning and alignment will take place early next week.  After this, the process will be repeated on the u/s set of legs.  We’ll then be ready for the PRY plates.
The designers, Steve Plate and Jason Tarrant, have been in the Hall discussing, supervising and fine-tuning the plans for the installation of the plates.  
At Keller, in New York, the plates for the south side have been completed.  The fit check has been “signed-off OK”.  Preparations are being made to air-ship the first set of plates.  The teams in the US and at Keller, as well as the local MICE and RAL logistics teams, are all primed to bring the plates into the UK and to the lab asap.  Unfortunately, as soon as the shipper picks up the plates to take them to the airport, we loose a degree of control.  Roy, Andy and Alan (Bross) will be watching closely to minimise any potential for delay that may arise from border-agency or customs bureaucracy.
Steady progress in Rack Room 2 is bringing the internal wiring towards completion.  So, up to the end of last week, the construction project remains on track.

Cooling demonstration:

In parallel to the construction work, the MIPO is focused on defining the interfaces and envelopes for the cooling-demonstration configuration.  It is important that we use all available engineering effort effectively.  The interfaces and envelope specifications must be completed to allow the design, fabrication and assembly work to continue in parallel to the completion of Step IV.


The run-up of ISIS has not gone quite according to plan.  The machine was not ready for MICE to take beam today (08Mar15).  Reluctantly, on Friday 06Mar15, Steve Boyd was forced to cancel the running planned for 08Mar15.  
However, in preparation for the 08Mar15 run, a second mock-data run was carried out on the 05Mar15.  This was a success; all systems required for running on 08Mar15 were commissioned.  While disappointing, the preparations will put us in a good position to begin operations on the weekend of 21Mar15.

The near future:

MICE Hall:  

The installation of the components of the PRY remains the priority for work in the Hall.  The Hall team is now laying plans to ensure the best use of the additional effort brought in to allow two-shift construction from 16Mar15.  This effort will be turned onto the installation of the south PRY plates as soon as they arrive on site.

Weekend operations from 21Mar15: 

The cancellation of the 08Mar15 run means that the final commissioning with beam of the equipment required to take data for TOF calibration will need to take place on the 21Mar15.  In the coming week the Operations team will block-out the activities that will take place during the present ISIS User Run (2014(03)).  
Shift sign-up for the weekend running is good, but not yet complete: please volunteer!  Shift-training will be organised on Fridays in advance of the weekend running.  
Looking ahead; we are planning to move towards full time running from June 2015 (ISIS User Run 2015(01).  It has been decided that operations pause for a scheduled maintenance day each Wednesday.  During this period, Wednesdays will also be exploited for training.
Operations meetings:
The second MICE/ISIS Operations Meeting will take place at 14:00 in the 16Mar15.  Steve Boyd will be in touch shortly about the essential preparations for this meeting and the agenda.  After this meeting, the regular weekly pattern of MICE Operations meetings will begin.  As you know, the pattern we have agreed, at least for User Run 2014(03), is:
  • Friday’s 11:00: In person Operations Meeting (modelled on the current Hall meeting).  All those involved in the weekend’s running are required to attend in person.  There will be no phone connection;
  • Monday’s 15:00: A meeting to debrief on the weekend’s operation and to plan the coming week.  Those involved in the preceding weekend’s run are expected to attend in person.  Others with responsibility for equipment, operations or involved in running the following weekend are expected to attend.  A phone connection will be provided for remote participation.

Also available in: Atom